# Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction

### v1

@@ 4.2 Baselines
We use two baselines: published scores from prior studies and previous SOTA models run on the fused dataset. Regular baselines include classic attention-based methods (e.g., CMLA Wang et al. (2017)), pipeline approaches (Peng-Two-stage Peng et al. (2020)), sequence-to-sequence frameworks (BARTABSA Yan et al. (2021)), sequence-tagging methods (GTS Wu et al. (2020), JET-BERT Xu et al. (2020)), MRC-based models (COM-MRC Zhai et al. (2022), Triple-MRC Zou et al. (2024)), graph-based approaches (EMC-GCN Chen et al. (2022), DGCNAP Li et al. (2023)), and the SOTA MiniConGTS Sun et al. (2024).
In addition to baselines on separate datasets, we selected two state-of-the-art models with publicly available and runnable code: the most recent model, MiniConGTS Sun et al. (2024), and the earlier state-of-the-art for pair extraction, BARTABSA Yan et al. (2021). We used their original configurations and data processing pipelines, modifying only the training data to include all four training sets from 14lap, 14res, 15res, and 16res. The models were then separately evaluated on the four corresponding test sets.
Models 14Res 14Lap 15res 16res
P R F1 P R F1 P R F1 P R F1
@@
SOTAs on fused dataset
BARTABSA-fused 75.40 76.76 76.07 72.36 63.40 67.59 88.85 92.77 90.77 87.06 87.32 87.19
MiniConGTS-fused 81.06 78.82 79.92 75.62 71.61 73.56 93.63 92.72 93.17 78.80 81.82 80.28
Ours on fused dataset
Trans-AOPE ($w_{1}$:$w_{2}$ = 1:1) 91.88 80.83 86.00 85.87 74.68 79.89 94.88 88.25 91.45 94.63 89.10 91.80
Trans-AOPE ($w_{1}$:$w_{2}$ = 1:0) 93.80 80.39 86.58 88.60 74.26 80.80 95.07 88.95 91.91 93.10 89.01 91.01
Table 3: Comparison of different models on multiple datasets for AOPE task. Recall and precision values are omitted where they are not provided. Former best scores are underlined, and current best scores are bold. $\diamond$ are retrieved from Yan et al., 2021. $\bullet$ is retrieved from Zhao et al., 2020, and $\star$ are retrieved from Sun et al., 2024
Models 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
@@
SOTAs on fused dataset
BARTABSA-fused 71.05 72.33 71.68 63.66 56.01 59.59 88.98 92.77 90.77 83.82 84.07 83.95
MiniConGTS-fused 73.89 72.84 73.36 68.26 57.42 62.37 90.64 89.32 89.98 72.73 73.21 72.97
@@
Trans-ASTE ($w_{1}$:$w_{2}$ = 1:1) 89.30 77.10 82.76 83.81 69.64 76.07 93.80 86.71 90.11 92.18 86.81 89.42
Trans-ASTE ($w_{1}$:$w_{2}$ = 1:0) 91.70 78.24 84.44 84.63 70.59 76.98 94.31 88.11 91.11 92.78 87.58 90.11
Table 4: Comparison of different models on multiple datasets for ASTE task. Former best scores are underlined, and current best scores are bold. $\star$ are retrieved from Sun et al., 2024, and $\diamond$ is retrieved from Yan et al., 2021.
4.3 Results and Analysis
4.3.1 Study on Augmented Optimization
To assess the impact of weight compositions in Equation 10, we tested $(w_{1},w_{2})\in\{(1:0),(1:1),(1:10),(0:1),(10:1)\}$, where $w_{1}$ and $w_{2}$ are the weights for base and contrastive-based optimization, respectively, using a batch size of 4. Results for AOPE and ASTE are presented in Figures 5 and 5. For AOPE, contrastive-heavy weights ($w_{1}:w_{2}=1:1$ and $w_{1}:w_{2}=1:10$) performed well during early epochs (1–50), demonstrating the early advantages of contrastive optimization. After epoch 50, base ($w_{1}:w_{2}=1:0$) and balanced ($w_{1}:w_{2}=1:1$) optimizations provided greater stability and better final performance. In ASTE, base and balanced optimizations consistently outperformed others, indicating that ASTE primarily benefits from foundational learning, with limited advantages from contrastive optimization.
Similar trends were observed across other datasets, as shown in Appendix A.4, highlighting the broader applicability of these findings. Although contrastive-enhanced optimization does not deliver the best overall performance, it effectively improves efficiency in pair extraction tasks.
4.3.2 Comparison with Baselines
On AOPE
[…]
@@ 4.3.2 Comparison with Baselines
Results for ASTE are listed in Table 4. For the 14Res dataset, Trans-ASTE achieves the highest F1 score (84.44%), representing an APG of 12.76% over BARTABSA-fused. Similarly, it outperforms MiniConGTS-fused with an APG of 11.08%. In the 14Lap dataset, Trans-ASTE) achieves an APG value of 17.39% in F1 compared to BARTABSA-fused, and 14.61% over MiniConGTS-fused. For the 15Res dataset, Trans-ASTE) maintains a high level of performance, with an F1 score (91.11%) slightly leading both baselines. The overall F1 improvement is 0.34% compared to BARTABSA-fused and 1.13% compared to MiniConGTS-fused. The 16Res dataset also sees consistently high performance from Trans-ASTE), with the highest F1 score (90.11%), representing an APG of 6.16% over BARTABSA-fused.
Overall, combining datasets for all-for-one training does not always result in performance improvements across all datasets for SOTA models. However, when data is fused for our model, it consistently achieves performance gains across datasets, with significant improvements observed in 14Res, 14Lap, and 16Res. That said, the slightly lower performance on the 15Res AOPE task highlights potential areas for further optimization in specific scenarios.
5 Conclusion
@@
In this paper, we present an efficient transition pipeline for aspect-opinion pair extraction with linear time complexity $O(n)$, enhanced by a contrastive learning-based optimization method. This approach eliminates the need to directly identify and extract individual tokens, avoiding token-level bias. It can be trained on datasets with the most comprehensive coverage of the designed actions, resulting in significant performance improvements across various datasets. Specifically, training our model on a well-covered fused dataset enables it to learn robust action combination patterns, leading to superior performance on all datasets. Our model surpasses SOTA models on the same fused dataset, setting new benchmarks for AOPE and ASTE tasks. This approach not only enhances model efficiency but also demonstrates its effectiveness and versatility in improving performance.
@@
The primary limitation of the transition-based pipeline lies in its reliance on relatively large and diverse datasets to effectively learn action patterns. As evidenced by the findings in Figure 3, the model achieved better performance on the 15res test set when trained on datasets like 14res and 16res, which appear to provide richer and more diverse action patterns compared to the 15res training set. This observation highlights the inadequacy of relying solely on a single dataset, as it may lack the breadth and complexity needed to capture the full spectrum of action patterns required for robust model generalization. Without access to such diverse training data, the model’s ability to adapt to nuanced patterns or unseen contexts in the test data may be significantly compromised, emphasizing the importance of training on a fused or broader dataset to enhance its overall effectiveness.
Appendix A Appendix
A.1 Alternative Solution with State Changes
Since stack can hold more than two tokens, "is" can be moved to stack ($\beta_{3}\rightarrow\sigma_{3}$) before "Gourmet" and "food" merge ($[\sigma_{1},\sigma_{2},\sigma_{3}]\rightarrow[\sigma_{1\&2},\sigma_{3}]$). The detailed process is present in Table 5.
Phrase Action Stack ($\sigma$) Buffer ($\beta$) Aspect Opinion Pair
[] [$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
1 $SF$ [$\sigma_{1}$] [$\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
2 $SF$ [$\sigma_{1}$, $\sigma_{2}$] [$\beta_{3}$, $\beta_{4}$]
3 $SF$ [$\sigma_{1}$, $\sigma_{2}$, $\beta_{3}$] [$\beta_{4}$]
4 $M$ [$\sigma_{1\&2}$, $\beta_{3}$] [$\beta_{4}$]
5 $R_{n}$ [$\sigma_{1\&2}$] [$\beta_{4}$]
6 $SF$ [$\sigma_{1\&2}$, $\sigma_{4}$] []
A.2 Algorithm for Transition Actions
In Algorithm 1, while the core of the algorithm focuses on the action operations, the configuration also plays a vital role in its functionality.
The algorithm begins with a buffer, a list of input tokens representing the text to be parsed. These tokens are consumed during the parsing process and form the basis for token pair relationships. A critical component of the algorithm is the pair2sent mapping, which associates token pairs with their corresponding sentiment labels, allowing the algorithm to enrich the output with sentiment information, even for token pairs not explicitly provided in the target data (defaulting to a neutral label). Then, each parsing action (e.g., Shift, Left Relation, Right Relation) is mapped to a unique numeric ID. This mapping enables a compact representation of the sequence of parsing actions, which is useful for downstream tasks like training or evaluation. Throughout the parsing process, the algorithm records intermediate states (stack, buffer, actions, and sentiment labels) to maintain a complete history of its operations. Simultaneously, valid token pairs are stored in pairs, capturing the structured relationships in the text.
Algorithm 1 Argument-Opinion Pair Parsing Algorithm
1:buffer (list of input tokens), pair2sent (mapping of token pairs to sentiment labels), action2id (mapping of parsing actions to numeric IDs)
2:states (list of recorded parsing states), pairs (list of valid token pairs)
3:Initialize empty stack, actions, states, and pairs. $\triangleright$ Main processing loop: parse tokens from the buffer
4:while buffer is not empty do
5:     if len(stack) < 2 then
6:         stack.append(buffer.pop(0)) $\triangleright$ Apply $SH$: Shift
7:         Append $SF$ to actions.
8:         continue
9:     end if $\triangleright$ Form a pair from the top two elements in stack
10:     stackPair $\leftarrow$ (stack[-2], stack[-1])
11:     currentAction $\leftarrow$ getAction(stackPair, targetPairs)
12:     Append currentAction to actions. $\triangleright$ Handle relation formation actions ($LR$ and $RR$)
13:     if currentAction == $LR$ then
14:         Add stackPair to pairs.
15:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
16:         Apply stack[-2] and stack[-1] form a left-relation pair.
17:     else if currentAction == $RR$ then
18:         Add stackPair to pairs.
19:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
20:         Apply stack[-1] and stack[-2] form a right-relation pair.
21:     else
22:         sentimentLabel $\leftarrow 3$triangleright Default sentiment label.
23:     end if $\triangleright$ Record the current parsing state
24:     Record (stack, buffer, actions, sentimentLabel) in states. $\triangleright$ Modify the stack based on currentAction
25:     if currentAction == $L_{n}$ then
26:         Remove stack[-2] $\triangleright$ Apply $L_{n}$: Left Constituent Removal
27:     else if currentAction == $R_{n}$ then
28:         Remove stack[-1] $\triangleright$ Apply $R_{n}$: Right Constituent Removal
29:     else if currentAction == $M$ then
30:         Merge stack[-2] and stack[-1] $\triangleright$ Apply $M$: Merge
31:     end if
32:end while $\triangleright$ Finalize by reducing the remaining elements in the stack
33:while len(stack) > 1 do
34:     Remove the top element from stack $\triangleright$ Apply $ST$: Stop
35:end while $\triangleright$ Convert actions to numeric IDs for output
36:Map each action in states to its ID using action2id. return states, pairs.
A.3 Action Coverage on Different Datasets
14lap 14res 15res 16res fused
train (Precision) 99.99 99.99 99.99 99.99 99.99
train (Recall) 85.51 86.63 92.98 93.04 88.73
train (F1) 92.19 92.84 96.36 96.40 94.03
dev (Precision) 99.99 99.99 99.99 99.99 99.99
dev (Recall) 85.97 86.89 94.92 93.04 89.88
dev (F1) 92.46 92.99 97.39 96.40 94.67
test (Precision) 99.99 99.99 99.99 99.99 99.99
test (Recall) 86.66 87.54 91.47 92.42 89.41
test (F1) 92.85 93.36 95.54 96.06 94.41
total (Precision) 99.99 99.99 99.99 99.99 99.99
total (Recall) 86.02 87.03 92.56 92.79 89.00
total (F1) 92.48 93.06 96.14 96.26 94.18
Table 6: Transposed Transitional Action Coverage for 14Lap, 14Res, 15Res, 16Res, and fused Datasets.
The coverage of actions, as shown in Table 6, reveals some differences in the recall values across data sets. The 14Lap and 14Res datasets exhibit lower recall compared to the 15Res and 16Res datasets. This trend is observed across training, development, and testing phases, with recall values for 14 datasets remaining around 85-87%, while 15Res and 16Res achieve significantly higher recalls, exceeding 91% in most cases. The fused dataset strikes a balance with intermediate recall values, reflecting the integration of characteristics from all datasets. This highlights a relative weakness in capturing relevant actions in 14 datasets compared to 15 and 16 datasets.
A.4 Growth with Different Optimization Weight
Figures 7 and 7, 9 and 9, and 11 and 11 depict the performance growth with varying weights applied to the 14res, 15res, and 16res datasets, respectively. These figures highlight the influence of weight adjustments on model accuracy and efficiency in ASTE tasks. By examining the trends across these datasets, we demonstrate the early benefits of augmented optimization in enhancing performance.
Figure 6: Growth on different weights for AOPE task on 14res testing set.
Figure 7: Growth on different weights for ASTE task on 14res testing set.
Figure 8: Growth on different weights for AOPE task on 15res testing set.
Figure 9: Growth on different weights for ASTE task on 15res testing set.
Figure 10: Growth on different weights for AOPE task on 16res testing set.
Figure 11: Growth on different weights for ASTE task on 16res testing set.
A.5 Growth Comparison for AOPE and ASTE
The growth comparison for AOPE and ASTE is illustrated in Figures 13, 13, 15, and 15, corresponding to the 14res, 14lap, 15res, and 16res datasets, respectively. While ASTE generally follows the trends observed in AOPE, the fluctuations in ASTE are more extreme. These figures emphasize the sensitivity of ASTE to variations in model weights or dataset conditions, which results in sharper peaks and deeper troughs compared to the relatively smoother patterns seen in AOPE. This highlights the unique challenges posed by the ASTE task across different datasets.
Figure 12: Growth Comparison for AOPE and ASTE on 14res Testing set.
Figure 13: Growth Comparison for AOPE and ASTE on 14lap Testing set.
Figure 14: Growth Comparison for AOPE and ASTE on 15res Testing set.
Figure 15: Growth Comparison for AOPE and ASTE on 16res Testing set.

### v2

@@ 4.3 Baselines
We compare our models with two groups of baselines: (1) published results from prior studies that do not have runnable code available and (2) previous State-of-the-Art (SOTA) models re-trained by us using publicly available code base.
Group (1), as summarized in Table 1, include CMLA Wang et al. (2017), Peng-Two-stage Peng et al. (2020), BARTABSA Yan et al. (2021), GTS Wu et al. (2020), JET-BERT Xu et al. (2020), COM-MRC Zhai et al. (2022), Triple-MRC Zou et al. (2024), EMC-GCN Chen et al. (2022), DGCNAP Li et al. (2023), and the latest MiniConGTS Sun et al. (2024).
For re-trained models in group (2), we use two state-of-the-art models with publicly available and runnable code: the most recent model, MiniConGTS Sun et al. (2024) [^1] https://github.com/qiaosun22/MiniConGTS, which achieves SOTA performance on both ASTE and AOPE tasks with a complexity of $O(n^{2})$ (fixed constants are omitted for simplicity), and an earlier study, BARTABSA Yan et al. (2021) [^2] https://github.com/yhcc/BARTABSA, designed for the AOPE task with a complexity of $O(n)$. We use their original configurations and data processing pipelines, modifying only the training data to include all four training sets from 14lap, 14res, 15res, and 16res, when necessary. The models are then separately evaluated on the four test sets.
14res 14lap 15res 16res
In-domain setting P R F1 P R F1 P R F1 P R F1
@@
Trans-AOPE (Ours) 71.04 67.03 68.98 60.32 44.08 50.94 72.91 66.30 69.45 73.56 64.58 68.78
Combined-train setting (Training Sets)
MiniConGTS (14lap & 14res) 75.72 78.20 76.94 71.05 68.64 69.83 63.30 69.90 66.44 68.13 72.54 70.27
MiniConGTS (14res, 15res, & 16res) 78.68 76.78 77.72 56.86 48.31 52.23 94.54 92.48 93.50 77.65 75.22 76.42
MiniConGTS (14res, 14lap, 15res, & 16res) 78.22 77.11 77.66 76.32 64.19 69.74 93.35 91.99 92.67 76.27 76.79 76.53
BARTABSA (14res, 14lap, 15res & 16res) 75.40 76.76 76.07 72.36 63.40 67.59 93.56 94.43 93.56 87.06 87.32 87.19
Ours
Trans-AOPE (14lap & 14res) 91.95 83.24 87.38 90.60 79.88 84.91 73.99 73.19 73.59 74.84 69.94 72.31
Trans-AOPE (14res, 15res & 16res) 74.34 62.29 67.78 91.03 78.11 84.07 95.66 88.04 91.69 90.63 80.65 85.35
Trans-AOPE (14res, 14lap, 15res & 16res) 92.92 83.61 88.02 92.20 80.47 85.94 96.17 90.94 93.48 93.75 84.82 89.06
Table 5: Comparison of different models on multiple datasets for AOPE task. Recall and precision values are omitted where they are not reported. The former best scores are underlined, and current best scores are bold. Highlights are used for analysis. $\diamond$ are retrieved from Yan et al., 2021. $\bullet$ is retrieved from Zhao et al., 2020, and $\star$ are retrieved from Sun et al., 2024
5 Results and Analysis
In this section, we first investigate the effect of adding contrastive loss on performance to determine the optimal training loss configuration (section 5.1). Next, we present the results of experiments under different training data configurations, and compare our model with previous methods (section 5.2). We analyze our results in section 5.3.
5.1 Effect of Contrastive Loss
To investigate the impact of different weight configurations between the base loss and contrastive loss in Equation 9, we evaluated five ratios in $w_{base}:w_{con}$—–namely $1:0$, $1:1$, $1:10$, $0:1$, and $10:1$—–with a batch size of 4. The results on 14lap dataset for the AOPE task are shown in Figure 3. Notably, the contrastive-heavy settings where $w_{base}:w_{con}=1:1~{}\text{and}~{}1:10$ achieved stronger early-stage performance, indicating the initial advantages of adding contrastive optimization.
Figure 3: F1 score as a function of training epochs in the combined-train condition for the AOPE task on the 14lap test set, with various loss weight configurations. w1=base loss; w2=contrastive loss.
A similar pattern emerged in the ASTE task, where the base and balanced configurations consistently outperformed the other weight settings, and the balanced option exhibited faster improvements than the base setting. These trends were also observed across additional datasets. Consequently, all subsequent experiments used the balanced configuration $w_{base}:w_{con}=1:1$ for optimization.
5.2 Main Results
On AOPE
First, in the in-domain setting, presented in the upper part of Table 5, when models are trained on one dataset only, we observe that our Trans-model (Trans-AOPE) performs worse than previous models. [^3] Note that the numbers for previous models are retrieved from respective papers that describe the system, and we are assuming they were trained with only in-domain training set, since many of them did not report which training set is used.
Second, for the combined-train condition, Trans-AOPE has an advantage over MiniConGTS in all combinations of training sets. What is particularly striking is that, when trained with four training sets combined, Trans-AOPE achieves roughly 20 points or more increase in F1 on all test sets (from 68.98 to 88.02 for 14res, 50.94 to 85.94 for 14lap). This is not the case for MiniConGTS, which sees performance drop in two test sets (14res and 14lap), and negligible increase in one (16res), large increase in another (15res: 73.87 to 92.67, but still lower than Trans-AOPE, 93.48). This suggests that MiniConGTS is not adaptive and cannot reliably make use of a mixture of training data from two domains and multiple datasets. Similar trends are observed for BARTABSA.
What is more striking is that adding restaurant data to the training set (14res and 14lap) improves performance of Trans-AOPE on the laptop test set, from 50.94 to 84.91, but not for MiniConGTS, which drops from 73.23 to 69.83. Moreover, when trained only on the three restaurant datasets (14res, 15res, and 16res), Trans-AOPE achieves performance comparable to what it attains when the 14lap dataset is included in the training set (84.07 in the yellow row, compared to 84.91 in the blue row). In contrast, MiniConGTS, trained on those same three restaurant datasets, achieves an F1 score of only 52.23. This suggests that Trans-AOPE (unlike previous models) can reliably transfer what it has learned in one dataset to another, even if the data are from a different domain, in the AOPE task.
14res 14lap 15res 16Res
In-domain setting P R F1 P R F1 P R F1 P R F1
@@ 5.2 Main Results
Trans-ASTE (Ours) 65.46 59.38 62.27 53.44 41.42 46.67 66.28 61.96 64.04 64.82 59.23 61.90
Combined-train setting (Training Sets)
MiniConGTS (14lap & 14res) 72.11 74.48 73.28 62.50 60.38 61.42 56.48 62.38 59.28 64.57 68.75 66.59
MiniConGTS (14res, 15res, & 16res) 74.64 72.84 73.73 52.12 44.28 47.88 92.31 90.29 91.29 74.42 72.10 73.24
MiniConGTS (14res, 14lap, 15res & 16res) 73.89 72.84 73.36 68.26 57.42 62.37 90.64 89.32 89.98 72.73 73.21 72.97
BARTABSA (14res, 14lap, 15res & 16res) 71.05 72.33 71.68 63.66 56.01 59.59 91.30 92.99 92.13 83.82 84.07 83.95
@@
Trans-ASTE (14lap &14res) 88.05 80.51 84.11 88.11 74.56 80.77 69.55 67.03 68.27 64.53 62.80 63.65
Trans-ASTE (14res, 15res & 16res) 73.53 61.74 67.12 42.25 32.25 36.58 89.89 86.96 88.40 81.19 77.08 79.08
Trans-ASTE (14res, 14lap, 15res & 16res) 89.56 81.24 85.20 86.87 76.33 81.26 94.34 90.58 92.42 90.03 83.33 86.55
Table 6: Comparison of different models on multiple datasets for ASTE task. Recall and precision values are omitted where they are not reported. Highlights are used for analysis. The former best scores are underlined, and current best scores are bold. $\star$ are retrieved from Sun et al., 2024, and $\diamond$ is retrieved from Yan et al., 2021.
@@
From Table 6 we observe similar patterns as AOPE, where our Trans-model demonstrates worse performance in the in-domain setting but superior performance under the combined-train setting. In the in-domain setting, Trans-ASTE is on par with previous models in 15res, but behind MiniConGTS on the other three test sets. However, under the combined-train condition, we again observe that our Trans-model benefits from the mixture of training data. For instance, while testing on 14res, the performance gains a 20 points boost when adding the training data of 14lap to 14res (62.27 to 85.20). On the contrary, for MiniConGTS, we see a drop in performance from 75.59 to 73.28. For BARTABSA, there is a moderate improvement (65.25 to 71.68). When four training sets are combined, Trans-ASTE shows superior performance in all test sets except 15res, usually by a large margin (gray rows in Table 6). This is especially true for 14lap, in which Trans-ASTE is the only model benefiting tremendously from adding restaurant data (performance increases from 46.67 to 81.26).
The only pattern different from Trans-AOPE is that when training on three restaurant datasets and testing on 14lap (yellow row), Trans-ASTE has very low performance of 36.58 F1, much lower than the previous 84.07 F1 in Trans-AOPE. We believe that this is due to the inability of the sentiment tagging system to transfer from one domain to another, rather than the issue of the aspect–opinion pair extraction of our transition-based pipeline. We thus find that Trans-model demonstrates stronger transfer performance in the AOPE task compared to ASTE, possibly because AOPE only requires predicting actions to extract aspect–opinion pairs. In contrast, ASTE requires both identifying these pairs and assigning a sentiment tag to form triplets, and sentiment tags tend to be more domain-specific.
5.3 When and Why is Trans-model Better?
When: Trans-model is better with combined training sets and in multi-domain generalization. From the results above, it seems clear that the Trans-model in the two tasks are better when trained on multiple datasets combined, rather than one dataset alone. Our results also suggest that the added training data do not have to be in the same domain: when trained on 14lap & 14res and tested on either 14lap or 14res, the F1 score is at least about 10 percentage points better than trained on only one of the two datasets. It shows that our Trans-model is capable of learning from multiple domains, and seems to be able to transfer its knowledge from the laptop domain to the restaurant domain and vice versa, unlike MiniConGTS, which seems to be more sensitive to the domain of the data. However, further research is need to better understand the discrepancy between Trans-AOPE and Trans-ASTE models in cross-domain transfer, and improve the cross-domain transfer ability of the sentiment tagger.
Why: Trans-model can learn more actions in data from diverse domains. We believe that the proposed trans-models excel when trained on multiple datasets and show considerable generalization ability for two main reasons. First, by predicting actions instead of tokens, it avoids token-level biases and prevents overfitting on token-level patterns specific to restaurant datasets. Additionally, trans-models perform pair extraction after the aspect–opinion relationship is established, which enables the model to capture contextual relationships more effectively. These design decisions work together to significantly enhance the overall effectiveness and robustness of the trans-models.
6 Conclusion and Future Work
@@
In this paper, we present an efficient transition pipeline for the extraction of aspects-opinion pairs with linear time complexity $O(n)$, enhanced by a contrastive-based optimization method. This approach obviates the need to directly identify and extract individual tokens, thereby mitigating token-level bias. It can be trained on a combination of diverse datasets that offers the most comprehensive coverage of the actions needed, resulting in significant performance improvements across various datasets. Specifically, training our model on a well-covered fused dataset enables it to learn robust action patterns, leading to superior performance on all datasets. Our model surpasses retrained baseline models on the same fused dataset, establishing new state-of-the-art results for both AOPE and ASTE tasks.
As transition-based methods have remained relatively less explored in sentiment-related tasks, we believe our work shows a promising direction to employ such methods in aspect-based sentiment analysis. Future work can further examine the potential of transition-based models in other sentiment analysis tasks, as well as the generalization ability of these models in situations of multi-domain data. It is also important to better understand the cross-domain and multi-domain generalization ability of transition models, by experimentation on more domains, since only two domains are involved in this work.
@@
Although the Trans-model demonstrates robust generalization capability, its reliance on larger datasets to effectively learn action patterns remains a notable limitation of the transition-based pipeline. This issue is evident in our results: while it is not necessary to use the same dataset for both training and testing, the model performs better when trained on blended datasets rather than on a single, limited one. Consequently, if the training data lack sufficient action patterns, the model’s ability to handle nuanced or previously unseen contexts can be significantly compromised. These findings underscore the importance of training on a combined or broader dataset to enhance the model’s overall effectiveness.